OpenStudy (anonymous):

Using complete sentences, explain each step in simplifying the ratio (x^2+8x)/(x^2+10x+16) and provide the simplified form.

OpenStudy (anonymous):

Factor the top and bottom and cancels factors

OpenStudy (anonymous):

i know that my answer is (x)/(x+2)

OpenStudy (anonymous):

apply difference of two squares rule: a^2+b^2=(a-b)(a+b) I got this..

OpenStudy (anonymous):

a^2+b^2 does not equal to (a-b)(a+b)

OpenStudy (anonymous):

a^2 - b^2=(a-b)(a+b)

OpenStudy (anonymous):

that is what i meant. my apologies

OpenStudy (anonymous):